Flying to Maui with a tailwind a plane averaged 158km/h. On the return trip the plane only averaged 112km/h while flying back into the same wind. Find the speed of the wind and the speed of the plane in still air.
I don't feel they have given me enough information to solve this problem or am I missing something?

Answer by ptaylor(2198)

I BELIEVE THERE IS ENOUGH INFO. LETS SEE WHAT WE COME UP WITH
Let r=rate (speed) of the plane in still air
and let w=rate (speed) of the tail wind
Now we are told that
r+w=158 ------------------------eq1
and
r-w=112------------------------------eq2
Add eq1 and eq2 and we get:
2r=270 divide both sides by 2
r=135 km/hr-----------------------speed of plane in still air
r-w=112
135-w=112 subtract 135 from both sides
135-135-w=-135+112 collect like terms
-w=-23 km/hr mulitply both sides by -1
w=23 km/hr ------------------------speed of the wind
ck
r+w=158
135+23=158
158=158
and
r-w=112
135-23=112
112=112
Hope this helps---ptaylor